# Promoting Constructive Deliberation: Reframing for Receptiveness

### v2

@@ 5.5 Toxicity effects on perceived reception
These results suggest that the receptive reframes have higher impact on the more toxic comments. Neither paraphrase nor receptive baseline generations were perceived to be any more receptive on toxic comments, further drawing the distinction between our reframing strategies and the baselines.
@@ 7 Conclusion
We curate six linguistic strategies from social science literature that signal receptiveness to another person. There is overlap between some of these strategies, particularly agreement with acknowledgement and grounding. While the overlap with grounding was along expected lines (as discussed in Section 4), overlap with acknowledgement was unexpected. It appears that the agreement prompt does not always reframe replies with explicit agreement and instead appears to simply acknowledge the speaker’s perspective. This is likely because we prompt the model to make “explicit any agreement with the comment.” and in the absence of agreement expressed in the original reply, the model appears to default to acknowledge rather than to agree. For a more distinct strategy, we might instead study token agreement, where we don’t try to make any agreement explicit (because there usually isn’t any) and instead simply reframe in the form “yes, but…”.
@@ Acknowledgements
We thank the online workers who provided annotations for this study. This research was supported in part by the Knight Foundation and by Good Systems [^4] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. We acknowledge the Texas Advanced Computing Center (TACC) for compute resources. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ Appendix D Additional Qualitative Examples
Examples of receptiveness reframes for comment-reply pairs are listed in Table 5. Examples of predicted contradictions between original and reframed replies are shown in Table 6. Examples of original replies that might be seen as less reasonable than the model-generated reframed reply are shown in Table 7.
@@
Well creepy isnt just a nickname anymore. It is an earned title. And let the games begin. Like trump hasn’t done the same shit? Tunnel vision at work Are you saying that Trump hasn’t done things that could also be considered creepy? It sounds like you might not be considering everything.
whales are not fish any more than seals or otters are. is that what u meant? I am positive that is not what the artist meant. Saving anything doesn’t mean fuck everything else. I understand your point about whales not being fish. But I’m sure the artist intended to convey saving wildlife in general, not disregarding other species.
No, this is going to be the start of the end of the republican party. This idiot is a genius in campaigning. I am dooming pretty hard for 22 and 24. I also think this could be the start of the end for the Republican party, but I’m worried about ’22 and ’24 since this "idiot" is actually a genius in campaigning.

### v3

@@ 5.5 Toxicity effects on perceived reception
These results suggest that our receptive reframes have higher impact on the more toxic comments. Neither paraphrase nor receptive baseline generations were perceived to be any more receptive on toxic comments, further drawing the distinction between our reframing strategies and the baselines.
@@ 6 Discussion
Beyond receptiveness
The framework described in this paper could be easily applied to address factors beyond receptiveness, which is not the only deliberative ideal that helps promote constructive deliberation. To encourage quality discussion across disagreements, deliberation scholars have identified multiple deliberative ideals in prior work, such as rationality (reason-giving) and common good orientation Bächtiger et al. (2018). While our receptiveness reframing aims to modify the tone of the discussion while preserving the meaning, one could use the same framework of incorporating low-level linguistic strategies from social science theory, but focus on improving the substance of the discussion. For example, to encourage more rational discussions, one could rewrite comments to both ask and make (or nudge users to make) more justification of claims made. Similarly, one could aim to rewrite comments to make arguments with a common good orientation rather than ones with narrow group interests.
@@ 7 Conclusion
We curate six linguistic strategies from social science literature that signal receptiveness to another person. There is overlap between some of these strategies, particularly between agreement, acknowledgement, and grounding. As discussed in Section 4, the overlap between these strategies was along expected lines. It appears that the agreement prompt does not always reframe replies with explicit agreement and instead appears to simply acknowledge the speaker’s perspective. This is likely because we prompt the model to make “explicit any agreement with the comment.” and in the absence of agreement expressed in the original reply, the model appears to default to acknowledge rather than to agree. For a more distinct strategy, we might instead study token agreement, where we don’t try to make any agreement explicit (because there usually isn’t any) and instead simply reframe in the form “yes, but…”.
@@
Lack of comparison between strategies
While we would have liked to evaluate the effectiveness between pairs of strategies, the size of our dataset does not allow for meaningful tests of significance for multiple pairwise comparisons and interaction analyses. While we illustrate how all strategies show strong improvements over the baselines, further between-strategy analysis would be valuable.
30 word limit to comments and replies
As we mention in Section 3.2, 30-word comments and replies are around the average word length of all comments and replies in our dataset. However, we note that the effect of reframing on longer messages may be different.
@@ Acknowledgements
We thank the online workers who provided annotations for this study, and the anonymous reviewers for their valuable feedback. This research was supported in part by the Knight Foundation and by Good Systems [^4] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. We acknowledge the Texas Advanced Computing Center (TACC) for compute resources. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ Appendix D Additional Qualitative Examples
Examples of receptiveness reframes for comment-reply pairs are listed in Table 5. Examples of predicted contradictions between original and reframed replies are shown in Table 6. Examples of original replies that might be seen as less reasonable than the model-generated reframed reply are shown in Table 7. Examples of comments and replies for varying levels of reply toxicity are shown in Table 8.
@@
“Well creepy isnt just a nickname anymore. It is an earned title. And let the games begin.” “Like trump hasn’t done the same shit? Tunnel vision at work” “Are you saying that Trump hasn’t done things that could also be considered creepy? It sounds like you might not be considering everything.”
“whales are not fish any more than seals or otters are. is that what u meant?” “I am positive that is not what the artist meant. Saving anything doesn’t mean fuck everything else.” “I understand your point about whales not being fish. But I’m sure the artist intended to convey saving wildlife in general, not disregarding other species.”
“No, this is going to be the start of the end of the republican party.” “This idiot is a genius in campaigning. I am dooming pretty hard for 22 and 24.” “I also think this could be the start of the end for the Republican party, but I’m worried about ’22 and ’24 since this "idiot" is actually a genius in campaigning.”
@@
Comment Reply Reply toxicity
“"Education funding from property taxes". How else do they expect free public schools?” “I don’t think property taxes in a certain area directly benefit that neighborhood’s schools. It just all goes into the general fund, as far as I know.” 0.006
“The GOP sucks but are we not going to push back against representative Waters’ comments? She should not have said what she said… "we need to be more confrontational"” “What is wrong with that? She understands the people. If there was no murder Minnesota be burning right now” 0.251
“How do all of these people get shot when it’s almost impossible to own a gun in NYC” “It’s really not that hard to get a gun in New York… Maybe a long gun but you can kill people with a pistol.” 0.508
“We may disagree on stuff, but the left doesn’t "hate America.” “Bullshit they don’t. Oh, sure. They don’t hate the land. They just hate the founders, our culture, everything we stand for and our entire way of life.” 0.751
“HTF did anyone think he make a great President? He was a P.O.S. all his life..” “Mindless idiots that are too fucking stupid to know that they’re fucking mindless idiots, that’s who.” 0.968
Table 8: Examples of Debagreement comments and replies for varying levels of reply toxicity.